# A Closer Look at Spatial-Slice Features Learning for COVID-19 Detection

### v1

@@ Abstract
Conventional Computed Tomography (CT) imaging recognition faces two significant challenges: (1) There is often considerable variability in the resolution and size of each CT scan, necessitating strict requirements for the input size and adaptability of models. (2) CT-scan contains large number of out-of-distribution (OOD) slices. The crucial features may only be present in specific spatial regions and slices of the entire CT scan. How can we effectively figure out where these are located? To deal with this, we introduce an enhanced Spatial-Slice Feature Learning (SSFL++) framework specifically designed for CT scan. It aim to filter out a OOD data within whole CT scan, enabling our to select crucial spatial-slice for analysis by reducing 70% redundancy totally. Meanwhile, we proposed Kernel-Density-based slice Sampling (KDS) method to improve the stability when training and inference stage, therefore speeding up the rate of convergence and boosting performance. As a result, the experiments demonstrate the promising performance of our model using a simple EfficientNet-2D (E2D) model, even with only 1% of the training data. The efficacy of our approach has been validated on the COVID-19-CT-DB datasets provided by the DEF-AI-MIA workshop, in conjunction with CVPR 2024. Our source code will be made available.
@@ 1 Introduction
With the rapid development of deep learning (DL), DL methods [17, 51, 26, 18, 25, 48, 63] have gained prominence for their ability to quickly and accurately identify COVID-19 features while efficiently handling large volumes of data. Furthermore, convolutional neural networks (CNNs) have proven to be more effective than methods based on frequency-domain [68, 49] and low-level features for CT image analysis [41].
@@
Inspired by [57], Tran et al. presented that factorizing the 3D convolutional filters (R3D) into separate spatial and temporal components (R(2+1)D) can yielding significantly gains in accuracy for action recognition. Its effectiveness have been demonstrated by several works on the fields of Video Understanding (VU)[42, 20, 44, 6] and Human Action Recognition (HAR) [64, 58]. One video may contains huge redundant information, such as noise from the audio track or each frame, and meaningless background, these factors make it difficult to train the model well [7], resulting in a significant increase in potential costs for data collecting. Likewise, CT scans can be regarded as a special case of video, it also contains various noise resulted from machine aging, and non-important spatial-slice pattern due to its imaging process [53]. Therefore, the different convolution methods on CT-scan is worthy of discussion.
@@
In this work, we introduce a Spatial-Slice Feature Learning (SSFL++) method, an unsupervised approach designed to reduce computational complexity by effectively removing out-of-distribution (OOD) slices and redundant spatial information. Furthermore, previous works [30, 11] have struggled to identify the most influential slices while considering global sequence information. Based on this observation, we believe there is room for improvement. Therefore, we propose a Kernel-Density-based Slice Sampling (KDS), which leverages Kernel Density Estimation to simultaneously achieve both objectives. Experimental results have demonstrated our 2D model’s outstanding performance, even in the face of data insufficiency.
@@ 2.3 Spatiotemporal Feature Learning for Video
Video analysis is crucial for computer vision, as videos contain far more information than single images. This analysis focuses on extracting spatiotemporal features, with traditional methods relying on optical flow [55, 8] and trajectory analysis [50, 67]. With the advent of deep learning (DL), a strategy employing 2D Convolutional Neural Networks (CNNs) was proposed [22, 65]. This strategy includes temporal feature pooling to aggregate features from different frames for classification. Subsequently, approaches combining CNNs with Recurrent Neural Networks (RNNs) and Long Short-Term Memory (LSTM) networks [1] were introduced, aiming to capture the long-range dependencies across various frames. 3D convolution kernels (C3D [56], I3D [9]) are used in video understanding, capturing channel interactions and local interactions simultaneously. However, they lead to a computational burden and have been regarded as an inefficient approach.
@@
2D: 2D Convolution over the sampled slices.
The use of 2D convolutional networks for extracting spatio-temporal features from 3D-cube data faces certain limitations, such as the requirement for strong spectral band or temporal continuity. Without these prerequisites, 2D convolutions may struggle to perform effectively due to their focus on spatial features and a lack of comprehensive sequence modeling. In applications involving CT scans, 2D convolutions are generally considered less effective compared to architectures like 2+1D convolutions, CNN-LSTM, or CNN-RNN, which are capable of capturing spatio-temporal features more efficiently. However, previous 2D CNN approaches often involve preprocessing, where crucial slices are selected and sampled to serve as inputs for the network. This sampling process tends to be overly simplistic, for instance, by manually selecting slices with the least artifacts or best quality, or randomly selecting a few slices to train a 2D CNN model. This limits the network’s potential for global sequential modeling.
@@ 4.2 Density-aware Slice Sampling
$K(x,x^{\prime})=\exp\left(-\frac{\|x-x^{\prime}\|^{2}}{2\sigma^{2}}\right)$ (7)

### v2

@@ Abstract
Conventional Computed Tomography (CT) imaging recognition faces two significant challenges: (1) There is often considerable variability in the resolution and size of each CT scan, necessitating strict requirements for the input size and adaptability of models. (2) CT-scan contains large number of out-of-distribution (OOD) slices. The crucial features may only be present in specific spatial regions and slices of the entire CT scan. How can we effectively figure out where these are located? To deal with this, we introduce an enhanced Spatial-Slice Feature Learning (SSFL++) framework specifically designed for CT scan. It aims to filter out OOD data within the entire CT scan, enabling us to select crucial spatial slices for analysis by reducing 70% redundancy totally. Meanwhile, we proposed Kernel-Density-based slice Sampling (KDS) method to improve the stability during training and inference stage, therefore speeding up the rate of convergence and boosting performance. As a result, the experiments demonstrate the promising performance of our model using a simple EfficientNet-2D (E2D) model, even with only 1% of the training data. The efficacy of our approach has been validated on the COVID-19-CT-DB datasets provided by the DEF-AI-MIA workshop, in conjunction with CVPR 2024. Our code is available at https://github.com/ming053l/E2D.
@@ 1 Introduction
With the rapid development of deep learning (DL), DL methods [17, 51, 26, 18, 25, 48, 63] have gained prominence for their ability to quickly and accurately identify COVID-19 features while efficiently handling large volumes of data. Furthermore, convolution neural networks (CNNs) have proven to be more effective than methods based on frequency-domain [68, 49] and low-level features for CT image analysis [41].
@@
Inspired by [57], Tran et al. presented that factorizing the 3D convolution filters (R3D) into separate spatial and temporal components (R(2+1)D) can yielding significantly gains in accuracy for action recognition. Its effectiveness have been demonstrated by several works on the fields of Video Understanding (VU) [42, 20, 44, 6] and Human Action Recognition (HAR) [64, 58]. One video may contains huge redundant information, such as noise from the audio track or each frame, and meaningless background, these factors make it difficult to train the model well [7], resulting in a significant increase in potential costs for data collecting. Likewise, CT scans can be regarded as a special case of video, it also contains various noise resulted from machine aging, and non-important spatial-slice pattern due to its imaging process [53]. Therefore, the different convolution methods on CT-scan is worthy of discussion.
@@
In this work, we introduce a Spatial-Slice Feature Learning (SSFL++) method, an unsupervised approach designed to reduce computational complexity by effectively removing out-of-distribution (OOD) slices and redundant spatial information. Furthermore, previous works [30, 11] have struggled to identify the most influential slices while considering global sequence information. Based on this observation, we believe there is room for improvement. Therefore, we propose the Kernel-Density-based Slice Sampling (KDS) strategy, which leverages Kernel Density Estimation to simultaneously achieve both objectives. Experimental results have demonstrated our 2D model’s outstanding performance, even in the face of data insufficiency.
@@ 2.3 Spatiotemporal Feature Learning for Video
Video analysis is crucial for computer vision, as videos contain far more information than single images. This analysis focuses on extracting spatiotemporal features, with traditional methods relying on optical flow [55, 8] and trajectory analysis [50, 67]. With the advent of deep learning (DL), a strategy employing 2D Convolution Neural Networks (CNNs) was proposed [22, 65]. This strategy includes temporal feature pooling to aggregate features from different frames for classification. Subsequently, approaches combining CNNs with Recurrent Neural Networks (RNNs) and Long Short-Term Memory (LSTM) networks [1] were introduced, aiming to capture the long-range dependencies across various frames. 3D convolution kernels (C3D [56], I3D [9]) are used in video understanding, capturing channel interactions and local interactions simultaneously. However, they lead to a computational burden and have been regarded as an inefficient approach.
@@
2D: 2D Convolution over the sampled slices. The use of 2D convolution networks for extracting spatio-temporal features from 3D-cube data faces certain limitations, such as the requirement for strong spectral band or temporal continuity. Without these prerequisites, 2D convolutions may struggle to perform effectively due to their focus on spatial features and a lack of comprehensive sequence modeling. In applications involving CT scans, 2D convolutions are generally considered less effective compared to architectures like 2+1D convolutions, CNN-LSTM, or CNN-RNN, which are capable of capturing spatiotemporal features more efficiently. However, previous 2D CNN approaches often involve pre-processing, where crucial slices are selected and sampled to serve as inputs for the network. This sampling process tends to be overly simplistic, for instance, by manually selecting slices with the least artifacts or best quality, or randomly selecting a few slices to train a 2D CNN model. This limits the network’s potential for global sequential modeling.
@@ 4.2 Density-aware Slice Sampling
$K(x,x^{\prime})=\exp\left(-\frac{\|x-x^{\prime}\|^{2}}{2\sigma^{2}}\right),$ (7)